A sphere of ice cream is placed on your ice cream cone. Both have a diameter of 8 cm. The height of your cone is 12 cm. Will all the ice cream, fit into the cone if it melted? Show all data to support your conclusion.
If no, what would be the diameter of a sphere of ice cream that would fill the cone? (Nearest tenth of a centimeter)

Answer by Cromlix(4381)

Hi there,
Volume of a sphere
= 4/3 πr^3
= 4/3 π4^3
= 268.08 cm^3
Volume of cone
= 1/3πr^2h
= 1/3π4^2 12
= 201.06 cm^3
So melting ice cream would
over spill cone.
To fit in cone
4/3πr^3 = 201.06 cm^3
r^3 = 3(201.06)/4 (π)
r^3 = 48 cm
r = 3√48
r = 3.63
diameter = 7.27 cm
Hope this helps:-)